Katie is 2 years younger than her brother. In 12 years, her brother will be twice as old as she is now. How old is Katie now?

Found 2 solutions by ikleyn, advanced_Learner:
Answer by ikleyn(52788)

.
Katie is 2 years younger than her brother. In 12 years, her brother will be twice as old as she is now. How old is Katie now?
~~~~~~~~~~~~~~~~~~~~

B - K = 2            (1)   ("Katie is 2 years younger than her brother")
B + 12 = 2K          (2)   ("In 12 years . . . )

To solve this system, express B = K + 2 from (1) and then substitute it into 2).
You will get a single equation for K:

(K + 2) + 12 = 2K,  or

K + 14 = 2K,  or

K = 14. 

Answer.  Katie is 14 years now.

For age problems, see the lessons
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.


Answer by advanced_Learner(501)

katie age now x-2 12 years x-2+12
bro age now x 12 years x+12
x+12=2(x-2)
x=16
katie age now 14 12 years 26
bro age now 16 12 years 28
indeed 28=2*14